George invests 2490 dollars at a simple interest rate of 10.4 percent. How much is his investment worth after 27 months?
Answer #1

amount invested = A = $2490

interest rate = r = 10.4% = 0.104

period of investment = n= 27 months

period of investment (years) = t= n/12 = 27/12 = 2.25

worth of investment after 27 months = A*(1+(r*t)) = 2490*(1+(0.104*2.25)) = 2490*1.234 = $3072.66